sold a combined total of 211 sodas and hotdogs, number is soda was 45 morethan hotdogs.
Answer by checkley79(3039)

S+HD=211
S=HD+45
(HD+45)+HD=211
2HD+45=211
2HD=211-45
2HD=166
HD=166/2
HD=83 ANS.
S=83+45
S=128 ANS.
PROOF:
128+83=211
211=211